OpenStudy (anonymous):

I am given the coordinates of three points and am asked to determine if they form an isosceles triangle how would i do this?

OpenStudy (anonymous):

Take the distances between all three points. If two are equivalent, they are isosceles.

OpenStudy (anonymous):

In general: $$A(x_0, y_0) \\ B(x_1, y_1) \\ C(x_2, y_2) \\ s_0 = \sqrt{(x_1 - x_0)^2 + (y_1 - y_0)^2} \\ s_1 = \sqrt{(x_1 - x_2)^2 + (y_1 - y_2)^2} \\ s_2 = \sqrt{(x_2 - x_0)^2 + (y_2 - y_0)^2} $$ If \(s_x\) for any \(x \in \{0, 1, 2\}\) have two equivalencies, it is isosceles.

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.